Assume that the partial pressure of sulfur dioxide, PSO2, is equal to the partial pressure of dihydrogen sulfide, PH2S, and therefore PSO2=PH2S. If the vapor pressure of water is 28 torr , calculate the equilibrium partial pressure of SO2 (PSO2) in the system at 298 K.
Answer #1

SO2(g) + 2H2S (g)-------> 3S + 2H2O(g)

\DeltaG = \DeltaGf products -\DeltaGf reactants

        = -2*228.59+0 -( -371.08 +2*-33.05)

      = -20KJ/mole   = -20000J/mole

\DeltaG = -RTlnKp

-20000 = -8.314*298*2.303logkP

logKp    = 20000/5705.8483

              = 3.5051

      Kp    = 103.5051    = 3199.6

PH2O = 28torr = 28/760 = 0.0368atm

Kp   = P2H2O/PSo2 P2H2S

3199.6   = (0.0368)2/x*x2

x3         = 0.0368*0.0368/3199.6   = 4.23*10-7

x          = 7.5*10-3 atm

PSo2 = x = 7.5*10-3 atm